OpenStudy (jack1):

A single phase transmission line possesses an inductive reactance of 15 ohms. Its supplied by an A/C source of 6000V (RMS). I need to know the voltage of at the end of the line if it has a capacitive load of 285 on it? Also, if you could help me work out the phase angle between the source voltage and the voltage across the capacitor? @pwnzerer

OpenStudy (kenljw):

apparently your considering resistance Vc = Vs (-j/2 pi 60 C)/(-j/2 pi 60 C + j2 pi 60 L) Vc = Vs (-j285/(-j285 + j15) Vc = Vs (-j285/(-j270) Vc = Vs 57/54 of course this is related to apparent power, all I used was the voltage divider rule with given impedances given